OpenStudy (anonymous):

\[Solve ~ for ~~~ 0 \le \phi \le 2 \pi\]\[3\sec^2 \phi - \sec \phi - 3 = 1\]

sam (.sam.):

3sec^2x-secx-2=0 (secx-1)(3secx+2)=0 secx=1, secx=-2/3 1/cosx =1 , 1/cosx=-2/3 cosx=1, cosx=-3/2

sam (.sam.):



OpenStudy (anonymous):

I don't think that's correct, because it would be -4 not -2

OpenStudy (anonymous):

But, please see if you can solve it! I need to go for lunch. I'll be back soon.

sam (.sam.):

x=0, 360

OpenStudy (anonymous):

That's not the answer in the book either... It says the answer is 0.72, pi, 5.56